# MIDAS: Multi-level Intent, Domain, And Slot Knowledge Distillation for Multi-turn NLU

### v2

@@ 3.3 Multi teachers knowledge distillation
Similarity loss $\mathcal{L}_{sim}$: The similarity loss is computed by maximising the logits similarities between the student model and teacher models. With this, the student model can learn the knowledge from the teacher models in the feature space, not only the prediction probabilities. The loss equation is:
$\mathcal{L}_{sim}=-\sum_{j}^{n_{T}}{\mathcal{F}_{sim}(v_{j},v_{s})},v_{j}=T_{j}(X),\,v_{s}=S(X)$ (1)
$\mathcal{L}_{tp}=\sum_{j}^{n_{T}}{cross\_entropy(v_{s},P_{j})},$
We experiment with diverse combinations of the aforementioned loss functions to assess their impact on the student’s performance across various datasets and NLU tasks. The summary of the loss function is described in Appendix B and the detailed result analysis can be found in Section 5.4.
4 Experimental setup
4.1 Datasets and Baselines
In this paper, we focus on multi-turn dialogue analysis in the dialogue state tracking (DST) domain, which consists of all three natural language understanding tasks, including intent classification, slot filling, and domain (topic) classification. Following by [9], we utilise two widely used benchmark datasets in multi-turn dialogue Natural Language Understanding (NLU): Multi-Domain Wizard-of-Oz 2.2 (MWOZ) and Machines Talking To Machines (M2M), especially used as benchmarks in the DST field. Details for each dataset are described in Appendix C.
@@ 4.1 Datasets and Baselines
Due to the limited number of baselines available for Multi-turn Dialogue Understanding, we initially adopted the three published results as baselines [2, 4, 9]. Additionally, we fine-tuned the pre-trained language models commonly used in the NLU domain. BERT-Base [^2] https://huggingface.co/bert-base-uncased is representative transformer encoder-based language model. RoBERTa-Base [^3] https://huggingface.co/roberta-base builds on BERT and adjusts hyperparameters by eliminating the next-sentence prediction objective. It also trains with larger mini-batches and learning rates. ALBERT-Base [^4] https://huggingface.co/albert-base-v2 is a BERT-based model that demonstrated superior performance by reducing the model size with factorised embedding parameterisation and cross-layer parameter sharing. SeqSeq [2] is an RNN with attention mechanisms, designed for the joint tasks of ID and SF. Slot-Gated [4] introduces a slot gate to capture the relationship between intent and slot, aiming to improve semantic understanding through global optimisation. Tri-level JNLU [9] is a pioneering model that incorporates domain information in the joint modelling of ID and SF.
4.2 Metrics and implementations
This paper evaluates the performance of baseline models and MIDAS in all three multi-turn dialogue tasks, including ID, SF, and DC for each dataset. Following by [2, 4, 9], the metrics for each task are shown as follows: Accuracy for ID and DC and F1 score for SF. Accuracy is the most commonly used metric for Intent Detection (ID) as determining the intent of an utterance is typically framed as a classification task. Accuracy is calculated as the ratio of correct predictions to the total number of tests. Domain Classification (DC) also employs accuracy as it is a classification task. On the other hand, Slot Filling (SF) employs F1 score. F1 score is directed towards assessing the prediction effectiveness for slot tokens. It computes an F1 score for each class and determines the token-based micro-averaged F1 score across all classes.
We introduce some implementation details in this section and the complete details in Appendix H. For Multi-teacher fine-tuning, we use BERT-Base, RoBERTa-Base and LLaMa2-7b [^5] https://huggingface.co/meta-LLaMA/LLaMA-2-7b as the teacher backbones and fine-tune them on each task. For fine-tuning LLaMa2-7b, we adopt an unmask strategy used in [35]. We use AdamW [36] and CrossEntropy loss to fine-tune the pre-trained models for 3 epochs. The learning rate is 5e-5 and is warm-uped linearly from 0 to 5e-5 during the first 10% training steps. The batch size is 32. For Multi-level Distillation, we use AdamW and the aforementioned losses to train the student with multi-level teachers. We use Squared Euclidean distance in algorithm 1 and cosine similarity in equation 1. For the combination of these losses, we sum them without any weight. We use the same optimiser, learning rate, warm-up strategy, and batch size as the one used in teacher fine-tuning, and use a vanilla Transformer encoder as a student.
5 Results
5.1 Overall performance
ID SF DC
MWOZ M2M MWOZ M2M MWOZ M2M
(ACC) (ACC) (F1) (F1) (ACC) (ACC)
BERT-Base 0.6534 0.8675 0.9218 0.8543 0.8667 0.8923
RoBERTa-Base 0.8424 0.9252 0.9748 0.9132 0.8675 0.8909
ALBERT-Base 0.6531 0.8654 0.9187 0.8542 0.8694 0.8919
SeqSeq 0.6641 0.9250 0.8543 0.9172 - -
Slot-Gated 0.6883 0.9327 0.8776 0.9279 - -
Tri-level JNLU 0.7849 0.9419 0.9798 0.9302 0.2572 0.8938
MIDAS (BERT) 0.8464 0.9427 0.9928 0.9856 0.8793 0.8952
MIDAS (RoBERTa) 0.8502 0.9377 0.9928 0.9813 0.8816 0.8945
We compare MIDAS with fine-tuned PLM baselines and published pioneering model results for two mainstream multi-turn natural language understanding tasks, Intent Detection and Slot Filling, with the same evaluation setup. Table 2 shows that MIDAS remarkably outperforms other baselines. To demonstrate the improvement achieved through MIDAS, we conduct experiments utilising two widely recognised multi-turn dialogue understanding datasets, MWOZ and M2M. Note that all baselines and MIDAS are fine-tuned for each of the two tasks individually. As detailed in Section 3.1, our approach involves the utilisation of pre-trained models, BERT or RoBERTa, for the fine-tuning of our three multi-level teacher models. These teachers encompass Intent Detection (ID), Slot Filling (SF), and Domain Classification (DC). It is important to highlight that MIDAS undergoes knowledge distillation from three distinct multi-level teachers, each specialising in sentence-level intent, word token-level slot, and conversation-level domain topic. Thus, Table 2 shows the results MIDAS (BERT) and MIDAS (RoBERTa) that all teachers are constructed using either the BERT or RoBERTa architecture. Two versions of MIDAS exhibit superior performance across both datasets, presenting outstanding outcomes with a slot-filling error rate below 2%. While the RoBERTa-Base model demonstrates superiority in MWOZ, the BERT-Base model excels in M2M. What should be noted is the performance difference between these models is not substantial, with both consistently outperforming other baseline models. In Intent Detection (ID) and Slot Filling (SF) tasks, MIDAS showcases notably higher performance compared to baselines. We also conduct experiments on the Domain Classification (DC) task with the same datasets to better compare the differences between MIDAS and other PLM baselines. However, while surpassing BERT-Base and ALBERT-Base, the performance difference is marginal. We assume that this discrepancy is attributed to the small number of the domain class. In contrast to other baseline models, Seq2Seq and Slot-Gated lack a structure incorporating domain information, making them unable to assess domain classification performance.
Overall, the observation highlights that bolstering multi-level conversation knowledge substantially improves the comprehension of each Natural Language Understanding (NLU) task. Specifically, enhancing results in intent detection is achievable by refining a student model through the distillation of multi-level knowledge, encompassing sentence-level intent, word-level slots, and conversation-level domain knowledge. The following two sections (Sections 5.2 and 5.3) delve into a more comprehensive exploration of multi-level teacher models and the combination of multi-level teachers.
5.2 Effect of pretrained model for teachers
Table 3-(a) shows the results of the effectiveness of using various pre-trained models as base models for all three multi-level teachers [^6] Note that the MIDAS (BERT) and MIDAS (RoBERTa) models are identical to those presented in the Table 2.. Compared to the high-achieving two encoder-based models, BERT and RoBERTa, the MIDAS (LLaMa) multi-level teachers produce lower performance [^7] Any decoder-only LM produces a similar low performance.. We assume a decoder-only model like LLaMa, primarily used for generating coherent and contextually relevant text, whereas BERT and RoBERTa are encoder-based models that have a deep understanding of context and relationships between words and excel in classification tasks. In addition to having multi-level teachers using a single pre-trained model, we adopt a mixed type of pre-trained model for preparing multi-level teachers. For instance, we can apply BERT as a pre-trained model for teaching sentence-level intent knowledge, utilise RoBERTa as a teacher model for word-level slots, and adopt LLaMa as a conversation-level domain topic teacher model. As shown in Table 3-(a), the result shows that using mixed types of pre-trained teacher models is less effective than employing a consistent single pre-trained model as the teacher. This implies that knowledge distillation from teachers with inconsistencies in their feature spaces may impede the learning process for a single student model.
ID SF DC
MWOZ M2M MWOZ M2M MWOZ M2M
(ACC) (ACC) (F1) (F1) (ACC) (ACC)
MIDAS (BERT) 0.8464 0.9427 0.9928 0.9850 0.8780 0.8952
MIDAS (RoBERTa) 0.8502 0.9377 0.9928 0.9813 0.8816 0.8945
MIDAS (LLaMa) 0.8403 0.9392 0.9912 0.9833 0.8702 0.8804
MIDAS (Mixed 1) 0.8472 0.9411 0.9839 0.9745 0.8808 0.8929
MIDAS (Mixed 2) 0.8473 0.9401 0.9928 0.9764 0.8769 0.8925
ID SF DC
MWOZ M2M MWOZ M2M MWOZ M2M
(ACC) (ACC) (F1) (F1) (ACC) (ACC)
ID-only 0.8406 0.9366 0.8590 0.9684 0.7977 0.7159
SF-only 0.8310 0.9377 0.9619 0.9718 0.2425 0.8930
DC-only 0.8408 0.9321 0.8888 0.9534 0.6330 0.8915
ID+SF 0.8422 0.9399 0.9924 0.9835 0.8760 0.8939
ID+DC 0.8400 0.9292 0.9923 0.9848 0.8756 0.8929
SF+DC 0.8376 0.9416 0.9923 0.9825 0.8760 0.8940
ID+SF+DC 0.8464 0.9427 0.9928 0.9850 0.8780 0.8952
(a)                                                                          (b)
Table 3: (a) The performance based on the type of teacher models. The MIDAS (BERT) and MIDAS (RoBERTa) are identical to those presented in the table 2 whose all teachers are either BERT or RoBERTa. MIDAS (LLaMa) refers to the outcome of utilising the LLaMa2-7b model as teacher models of all tasks. The MIDAS (Mixed 1 and 2) represents the outcome of mixed type teacher combination; Mixed 1: BERT (ID), LLaMa (SF) and RoBERTa (DC); Mixed 2: BERT (ID), RoBERTa (SF) and RoBERTa (DC). (b) The performance based on the type of teacher models. The first column indicates the type of teacher used. For example, ID+SF+DC uses all intent classification, slot filling, and domain classification teachers, while ID-only uses only the intent classification teacher. Only BERT is utilised as the teacher model. Results using RoBERTa are presented in Appendix D.
[…]
   MWOZ    M2M    MWOZ    M2M    MWOZ    M2M
(ACC) (ACC) (F1) (F1) (ACC) (ACC)
[…]
Table 4: The comparison of the diverse loss function combinations. Only BERT is utilised as the teacher model, and the results of RoBERTa are presented in Appendix E. The full names of each loss can be found in Section 3.3. We adopt two $\mathcal{L}_{KD}$ and $\mathcal{L}_{sce}$ as compulsory knowledge distillation loss and also explore three $\mathcal{L}_{rel}$, $\mathcal{L}_{sim}$, and $\mathcal{L}_{tp}$ for MIDAS. Scores in bold indicate leadership among the metrics, and underlined scores indicate the second-best.
5.3 Effect of combination of multi-level teachers
We explore the impact of incorporating each multi-level teacher (ID, SF, DC) in all three multi-turn dialogue understanding tasks. MIDAS is evaluated with individual teachers (ID, SF, DC), all possible pairs from {ID, SF, DC} [^8] Note that we do not adopt $\mathcal{L}_{rel}$ since it is not possible to adopt when there are two teachers., and then with all three teachers. Table 3 presents the results for each combination of teacher models for three different dialogue understanding tasks. Note that the table demonstrates the outcome of MIDAS (BERT) teachers, and we produce that of MIDAS (RoBERTa) in Appendix D. The experimental findings highlight that the ID+SF+DC combination attains the highest performance, underscoring the advantage of the student model integrating knowledge from all teachers for each task.
5.4 Effect of knowledge distillation loss function
As mentioned in Section 3.3, we conducted the loss function ablation study for MIDAS. This comprehensive evaluation aims to identify the most effective combinations that enhance the student model’s proficiency in handling different aspects of dialogue understanding across multiple NLU tasks. Note that we use $\mathcal{L}_{KD}$ and $\mathcal{L}_{sce}$ as compulsory knowledge distillation losses, and conduct an ablation study of three newly proposed multi-level teacher losses: $\mathcal{L}_{sim}$, $\mathcal{L}_{rel}$, and $\mathcal{L}_{tp}$. As shown in Table 4, the results indicate that incorporating $\mathcal{L}_{sim}$ with $\mathcal{L}_{rel}$ achieves the best or the second best performance across all tasks and datasets. Although $\mathcal{L}_{rel}$ and $\mathcal{L}_{sim}$ share a similar trend, their impact on model learning may be somewhat superior when employed independently, particularly when utilising $\mathcal{L}_{sim}$. While incorporating $\mathcal{L}_{tp}$ with the others led to a slight performance increase, it did not match the effectiveness observed with the sole application of the earlier losses. [^9] We conducted testing with $\mathcal{L}_{tp}$ only, it produces much lower performance than any others, which is expected.
5.5 Qualitative analysis: case study
Turn    Model                 Tokens (Slot)   Intent  Domain
1 Utterance boat, is, fine - -
GT B-RN, O, O affirm -
MIDAS (BERT) B-RN, O, O affirm -
BERT-Only O, O, O inform -
2 Utterance sure, ,, i, found, this, one, :, the, view - -
GT O, O, O, O, O, O, O, B-RN, I-RN offer -
MIDAS (BERT) O, O, O, O, O, O, O, B-RN, I-RN offer -
BERT-Only O, O, O, O, O, O, O, O, O offer -
3 Utterance how, about, the, ivy, or, boats, ? - -
GT O, O, B-RN, I-RN, O, B-RN, O select restaurant
MIDAS (BERT) O, O, B-RN, I-RN, O, B-RN, O select restaurant
BERT-Only O, O, B-RN, B-RN, O, B-RN, O select movie
Table 5: A Prediction example with a three-turn conversation on slot filling, intent detection, and domain classification results of each model. The green cell represents a result that matches the ground truth, the red cell indicates incorrect results, and the yellow cell indicates partially correct results. Additional case examples are articulated in Appendix F.
We further evaluate MIDAS with a qualitative assessment of the three NLU tasks on M2M. As shown in Table 5, we assume to have a three-utterance conversation [^10] The original conversation is longer but reduced in the visualisation due to page limitation., ‘boat is fine’, ‘sure, i found this one: the view’, ‘how about the ivy or boats?’. Based on the given conversation, we test all three NLU tasks, including intent classification, slot filling, and domain classification. The domain would be classified across the entire conversation, so we placed it only once to the end. We mainly compare with two models: 1) MIDAS (BERT), trained with three teachers $BERT_{ID}$, $BERT_{SF}$, and $BERT_{DC}$, and 2) BERT-only refers a single fine-tuned BERT (BERT-Base), which focus on only one task for each prediction.
Although the single fine-tuned BERT (BERT-only) can sometimes predict the slots correctly, it does not communicate/integrate with the word level and domain level classification. For example, (BERT-only) correctly tags B-RN (Restaurant Name) in the turn 3, while no-integration would produce a domain prediction as ‘movie’. Nevertheless, the student model exhibits the ability to identify (-RN)Restaurant Names even in the absence of domain labels when trained independently on the SF task and concurrently on the ID and DC tasks. Instances such as these validate our hypothesis that leveraging diverse knowledge levels from multi-turn conversation datasets can improve the understanding of individual natural language understanding tasks, outperforming the advantages of learning with single-level dialogue knowledge.
6 Conclusion and Limitation
This paper introduces a novel multi-level teacher knowledge distillation framework to enhance multi-turn natural language understanding (NLU). By fine-tuning pre-trained models at word, sentence, and document levels, we construct multi-level teachers, imparting their knowledge to a student model. Various loss functions are introduced and explored, and the experiment results demonstrate the framework’s effectiveness in improving the student model’s understanding across diverse NLU tasks. It shows better than the LLM result. However, there are some spaces for future work, including covering multilingual multi-turn dialogue. We believe this work will provide valuable insights into various aspects of dialogue knowledge for NLU and multi-level knowledge distillation.
Appendix A Related works
Table 6 presents a comparison of MIDAS with 23 previous joint NLU models. Recently, most NLU studies have embraced a joint learning model capable of handling all NLU tasks to mitigate error propagation inherent in pipelined approaches [39, 46, 8, 45]. The initial joint models employed traditional neural networks like RNN [2] and LSTM [5, 38, 41, 42, 43, 44] with attention mechanisms.
Only one previous study [9] conducted tests on domain classification. Hence, we chose [9] as a representative baseline. What sets the proposed model apart is its utilisation of multi-teacher knowledge distillation. While two previous works employed self-knowledge distillation and another two adopted one-teacher knowledge distillation, MIDAS represents the first attempt at employing multi-teacher knowledge distillation for joint learning in natural language understanding.
Appendix B Each losses of proposed loss function
The following losses are utilized to train the student model, each playing a distinct role:
  • $L_{kd}$: This loss facilitates the transfer of knowledge from the teacher models to the student model, enabling the student to mimic the general behavior of multiple teachers.
  • $L_{rel}$: This loss is designed to capture the relationships between different samples in the input data. It helps to align the student’s understanding with that of the teacher models and ensures a consistent perspective on the dataset.
  • $L_{sim}$: This loss encourages the student model to generate outputs that are similar to those of the teacher models in terms of their overall structure and distribution. It helps to maintain consistency between the student and teacher predictions.
  • $L_{sce}$: This loss function serves as the fundamental mechanism for training the student model. It entails the student learning to predict the correct labels associated with the input data.
  • $L_{tp}$: This loss leverages the predictions of the teacher models to provide additional supervision signals to the student model. It helps to guide the student towards making predictions that align with those of the teachers.
Appendix C Details of datasets
MWOZ [47] is specifically designed for Dialogue State Tracking (DST) and adopts the conventional human-vs-human Wizard of Oz approach across diverse domains, including attraction, bus, hospital, hotel, police, restaurant, taxi, and train. It incorporates 30 slot types and 11 intent types. The dataset comprises 8,437 dialogues, with an average of 5.68 turns per dialogue and 14.07 tokens per turn. Following by [2, 4, 9], we do not consider any multi-label samples but utilise the data with a single domain and intent.
M2M [48] is introduced with virtual agents and user-generated interactions to emulate goal-directed conversations through paraphrasing with templated utterances. M2M has movies and restaurant domains. The slots and intents are categorical, with 21 slot types and 15 intent types. The dataset comprises 1,500 dialogues, with an average of 9.86 turns per dialogue and 8.25 tokens per turn.
ID SF DC
MWOZ M2M MWOZ M2M MWOZ M2M
(ACC) (ACC) (F1) (F1) (ACC) (ACC)
ID-only 0.8339 0.8097 0.9079 0.9326 0.6183 0.7147
SF-only 0.8403 0.8945 0.9620 0.9434 0.2471 0.8917
DC-only 0.8469 0.8929 0.9547 0.9251 0.7521 0.8913
ID+SF 0.8451 0.9083 0.9928 0.9802 0.8734 0.8923
ID+DC 0.8373 0.9114 0.9921 0.9797 0.8763 0.8888
SF+DC 0.8453 0.9147 0.9927 0.9805 0.8707 0.8912
ID+SF+DC 0.8502 0.9377 0.9928 0.9813 0.8816 0.8945
Appendix D Combination-based ablation study
@@ 2 Utterance let, ’, s, go, with, the, view - -
Appendix E Loss function ablation study
[…]
   MWOZ    M2M    MWOZ    M2M    MWOZ    M2M
(ACC) (ACC) (F1) (F1) (ACC) (ACC)
[…]
We conducted the loss function ablation study for MIDAS with RoBERTa Teacher. This comprehensive evaluation aims to identify the most effective combinations that enhance the student model’s proficiency in handling different aspects of dialogue understanding across multiple NLU tasks. Note that we use $\mathcal{L}_{KD}$ and $\mathcal{L}_{sce}$ as compulsory knowledge distillation losses, and conduct an ablation study of three newly proposed multi-level teacher losses: $\mathcal{L}_{sim}$, $\mathcal{L}_{rel}$, and $\mathcal{L}_{tp}$. As shown in Table 9, the results indicate that incorporating $\mathcal{L}_{sim}$ with $\mathcal{L}_{rel}$ achieves the best or the second best performance across all tasks and datasets. Although $\mathcal{L}_{rel}$ and $\mathcal{L}_{sim}$ share a similar trend, their impact on model learning may be somewhat superior when employed independently, particularly when utilising $\mathcal{L}_{sim}$. While incorporating $\mathcal{L}_{tp}$ with the others led to a slight performance increase, it did not match the effectiveness observed with the sole application of the earlier losses.
Appendix F More case studies
We further evaluate MIDAS with a qualitative assessment of the three NLU tasks on MWOZ and M2M. In Table 8, we test all three NLU tasks, including intent classification, slot filling, and domain classification. The first two utterances are from M2M, while the rest are from MultiWOZ 2.2 (MWOZ). The first eight results come from MIDAS (BERT), trained with three teachers $BERT_{ID}$, $BERT_{SF}$, and $BERT_{DC}$, and BERT-only refers a single fine-tuned BERT (BERT-Base), whereas the remaining five results pertain to MIDAS (RoBERTa), trained with three teachers $RoBERTa_{ID}$, $RoBERTa_{SF}$, and $RoBERTa_{DC}$, and RoBERTa-only refers a single fine-tuned RoBERTa (RoBERTa-Base). Although the single fine-tuned BERT (BERT-only) or RoBERTa (RoBERTa-only) can sometimes predict the slots correctly, it does not communicate/integrate with the word level and domain level classification. Instances such as these validate our hypothesis that leveraging diverse knowledge levels from multi-turn conversation datasets can improve the understanding of individual natural language understanding tasks, outperforming the advantages of learning with single-level dialogue knowledge.
Appendix G Prompt method
G.1 Quantitative analysis

### v3

@@ Appendix A Related works
Appendix B Details of baselines
@@
Given the limited number of baselines available for Multi-turn Dialogue Understanding, we selected the following models as baselines. BERT-Base [^11] https://huggingface.co/bert-base-uncased is a transformer-based language model that serves as a standard benchmark. RoBERTa-Base [^12] https://huggingface.co/roberta-base improves upon BERT by removing the next-sentence prediction task and optimizing hyperparameters, such as using larger mini-batches and higher learning rates. ALBERT-Base [^13] https://huggingface.co/albert-base-v2 further enhances efficiency through factorized embedding parameterization and cross-layer parameter sharing, achieving better performance with fewer parameters. SeqSeq Liu and Lane (2016) is an RNN model with attention mechanisms, developed for joint intent detection ($ID$) and slot filling ($SF$) tasks. Slot-Gated Goo et al. (2018) introduces a slot-gating mechanism to capture the relationship between intent and slot labels, improving semantic understanding through global optimization. Lastly, Tri-level JNLU Weld et al. (2023) incorporates domain information for enhanced joint modeling of $ID$ and $SF$.
Appendix C Implementation details
C.1 Experiment hyperparameters
Hyper-parameter   Fine-tuning   Knowledge Distillation
Learning Rate 5e-5 5e-5
Batch Size 32 32
Warm-up Steps 10% of Max epoch 10% of Max epoch
Mex epoch 3 100
Stop Strategy Max Epoch Early Stopping on loss
Stop Patience - 10
Optimizer AdamW AdamW
Optimizer Weight Decay 1e-2 1e-2
Optimizer Betas 0.9, 0.999 0.9, 0.999
Margin in $\mathcal{L}_{rel}$ 0.2
Norm in $\mathcal{L}_{rel}$ 2
$\mathcal{F}_{D}$ in $\mathcal{L}_{rel}$ L2-Norm
Similarity in $\mathcal{L}_{sim}$ Cosine Similarity
Max Token Length 512 512
Table 10: The hyper-parameters used in our experiments.
C.2 Model details
C.3 Hardware information
     Model    Task   Learning Rate   Accuracy
M2M
LLaMa2-7b ID 0.0005 0.9121
0.00005 0.9392
0.000005 0.9093
SF 0.0005 0.9696
0.00005 0.9833
0.000005 0.9349
DC 0.0005 0.8895
0.00005 0.8804
0.000005 0.8375
Gemma-7b ID 0.0005 0.9204
0.00005 0.9357
0.000005 0.9102
SF 0.0005 0.9693
0.00005 0.9816
0.000005 0.9429
DC 0.0005 0.8799
0.00005 0.8840
0.000005 0.7890
MWOZ
LLaMa2-7b ID 0.0005 0.8021
0.00005 0.8403
0.000005 0.7952
SF 0.0005 0.9776
0.00005 0.9912
0.000005 0.9740
DC 0.0005 0.8411
0.00005 0.8702
0.000005 0.7026
  BERT   RoBERTa   LLaMa   Student
Architecture Encoder Encoder Decoder Encoder
Parameters 110M 125M 7B 58M
Layers 12 12 32 6
Heads 12 12 32 8
Hidden Dim. 768 768 4096 768
Feed Forward Dim. 3072 3072 11008 2048
Dropout Rate 0.1 0.1 0.0 0.3
Rank of LoRA - - 64
Alpha of LoRA - - 16
Dropout of LoRA - - 0.1
Table 12: The details of the models used in our work.
@@ 2 Utterance let, ’, s, go, with, the, view - -
Figure 3: Two examples for qualitative analysis: (a) shows the results on the M2M dataset, and (b) shows the results on the MWOZ dataset. Each example shows the results when MIDAS matches the ground truth. The three cells below each example display the type lists for slot, intent, and domain, and red text indicates errors in each column of the results table.
Appendix D In-depth PLM/LLM analysis
In this section, we provide an in-depth quantitative and qualitative analysis, incorporating a detailed comparison between PLMs and LLMs.
D.1 Compared with PLMs
We evaluate MIDAS with a qualitative assessment of the three NLU tasks on MWOZ and M2M, compared with two representative PLMs, BERT and RoBERTa. In Table 13, we test all three NLU tasks, including intent classification, slot filling, and domain classification. The first two utterances are from M2M, while the rest are from MultiWOZ 2.2 (MWOZ). The first eight results come from MIDAS (BERT), trained with three teachers $BERT_{ID}$, $BERT_{SF}$, and $BERT_{DC}$, and BERT-only refers a single fine-tuned BERT (BERT-Base), whereas the remaining five results pertain to MIDAS (RoBERTa), trained with three teachers $RoBERTa_{ID}$, $RoBERTa_{SF}$, and $RoBERTa_{DC}$, and RoBERTa-only refers a single fine-tuned RoBERTa (RoBERTa-Base). Although the single fine-tuned BERT (BERT-only) or RoBERTa (RoBERTa-only) can sometimes predict the slots correctly, it does not communicate/integrate with the word level and domain level classification. Instances such as these validate our hypothesis that leveraging diverse knowledge levels from multi-turn conversation datasets can improve the understanding of individual natural language understanding tasks, outperforming the advantages of learning with single-level dialogue knowledge.
D.2 Compared with LLMs
D.2.1 Quantitative analysis